# Bribe & Fork: Cheap PCN Bribing Attacks via Forking Threat

### v1

@@ 0.C.5 Discouraging Miners to Mine $txs_{1}$
Let us denote with $\sigma_{1}$ all strategies where player $N_{s}$ continues to mine the chain with $txs_{1}$ in $\mathcal{S}_{p_{1},f_{1}}^{T}$, and let $\sigma_{2}$ denote all strategies, where he forks $txs_{1}$ in $\mathcal{S}_{p_{1},f_{1}}^{T}$ and then continues to mine the newly created chain (we assume that the transaction $txs_{1}$ was not mined by the miner $N_{s}$, as when $f_{2}-f>\frac{f_{1}-f}{\lambda_{s}}$, then as shown in the proof of the Theorem 4, the player $N_{s}$ will choose to mine $txs_{u}$ in $\mathcal{S}_{p_{1}}^{T}$). It is easy to see that once the player $N_{s}$ successfully forks $txs_{1}$ and creates a new chain, all miners other than the miner that mined $txs_{1}$ will prefer to mine the newly created chain. As the miner who mined the transaction $txs_{1}$ has the mining power less than $\lambda_{s}$, we can say that once $txs_{1}$ is forked, the miners with the mining power at least $1-\lambda_{s}$ will continue mining the new chain. In summary, for some constant $C$, given similar arguments as in the previous proofs:
@@
The strategy $\sigma_{2}$ is dominating for the player $N_{s}$ who did not mine $txs_{p_{1}}$, and he will prefer to continue mining $txs_{p_{1}}$ whenever $f_{2}-f>\frac{2\overline{f}+2(f_{1}-f)}{\lambda_{s}}$. Note that the above analysis will be similar, even if some player other than the $N_{s}$ successfully starts a fork of $txs_{p_{1}}$.
@@
0.C.6 Encouraging the Strongest Miner to Use the Penalty Mechanism
[…]
Proof
By Theorems 4 and  4, utility of the player $N_{s}$ that chooses to create $txs_{p_{1}},txs_{p_{2}}$ and mine $txs_{p_{1}}$ in the first round [^4] In the analysis we omit the strategy where the player $N_{s}$ creates $txs_{p_{1}},txs_{p_{2}}$ and does not decide to mine $txs_{p_{1}}$. (strategy $\sigma_{p}$) is at least:
where $F^{\prime}_{1}=(\lambda_{s}(T-1)(f+B)+\lambda_{s}(R-T)(f+B)),$
Recall that we assume that all players with mining power less than $1\%$ have collective power at most $5\%$. As the players with mining power more than $0.01$ will prefer to mine $txs_{1}$ in the first place when $txs_{p_{1}}$ is not created, the utility of the player $N_{s}$ that does not decide to create $txs_{p_{1}}$ (strategy $\sigma_{1}$) is at most (by Lemma 3 and the Theorem 4):
On other hand the players may first decide to mine either $txs_{u}$ or $txs_{1}$ in the first round (strategy $\sigma_{1*}$). In the worst case scenario the block with $txs_{1}$ is not forked. What is more, whenever $1-\lambda_{p_{1}}$ miners decide to mine $txs_{u}$ in the first round, then all miners with mining power more than $0.01$ will make an attempt to mine $txs_{1}$. In conclusion, by Lemma 3 and the Theorem 4:
Which holds for any $f_{p_{1}}\geq f_{1}$ and $\lambda_{i}>0.05^{T/2}$.
Appendix 0.D Experimental Analysis
In this section we conduct the analysis of the fee structures on the the real-world data.
0.D.1 Fee Structure
We analyze the real-world data on fees earned by miners on Bitcoin. In particular, we explore the Bitcoin historical data from 2022 ([1]). In this period, the block reward was the sum of the block reward base, 6.25 BTC, and the fees for including transactions. The weekly average sum of the block fees ranged between 0.05 and 0.225 BTC as shown in Figure 3. The average fee of a single transaction fluctuated proportionally between $0.00004$ and $0.00016$ BTC as shown in Figure 4. We can thus fairly assume that any block typically contains on average $1500$ transactions, the sum of fees of an average block of transactions is around $0.15$ BTC and the average fee of a single transaction is approximately $10^{-4}$ BTC.
Figure 3: Average amount of fees aggregated in a block for each week in 2022 ([1]).
Figure 4: Average transaction fee of a single transaction for each week in 2022 ([1]).
0.D.2 Hash Power Analysis
Figure 5: Average mining hashrate week by week of the most significant Bitcoin miners in 2022 [1].
Device Hashrate (Hashes/second)
Nvidia GeForce RTX 4090 $2.75\cdot 10^{8}$
Nvidia GeForce RTX 3090 $9.4\cdot 10^{7}$
Bitmain Antminer S19 XP Hyd $2.55\cdot 10^{14}$
Baikal BK-G28 $2.8\cdot 10^{10}$
Table 2: Hash rates of commonly used Bitcoin mining devices [4].
The weakest miner. In the previous work [10], the authors assumed that the fraction of the mining power held by the weakest miners is approximately $10^{-3}$. The data depicted in the Table 2 and in the Figure $2(b)$ shows that it is a highly unrealistic assumption for Bitcoin. A very strong GPU which is often used by individual miners for mining has hashrate of around $3\cdot 10^{8}$ (Table 2), whereas the collective mining power in Bitcoin can be estimated to $2\cdot 10^{20}$ (Figure $2(b)$). In conclusion, we cannot expect that every individual miner has more than $10^{-12}$ probability of mining a block.
Strong miners. On the other side of the spectrum, top mining pools continuously maintain over $20\%$ of total hashpower. To calculate the share of the hashpower of a big mining pool in a given period, we count the number of the blocks mined by it. Then we divide this number by the total number of mined blocks in the period. Figure 6(a) and Figure 5 show that, in the most popular blockchains, most of the time, one of the miners (mining pools) has the collective power higher than $20\%$.
Pool Number of mined blocks Percentage of total mined blocks
Foundry USA $11851$ $22.281\%$
AntPool $8670$ $16.301\%$
F2Pool $7848$ $14.755\%$
Binance Pool $6164$ $11.589\%$
ViaBTC $5162$ $9.705\%$
Poolin $4438$ $8.344\%$
BTC.com $2252$ $4.234\%$
SlushPool $2142$ $4.027\%$
Luxor $1362$ $2.561\%$
SBI Crypto $963$ $1.811\%$
Braiins Pool $715$ $1.344\%$
unknown $698$ $1.312\%$
MARA Pool $307$ $0.577\%$
Terra Pool $149$ $0.280\%$
ULTIMUS POOL $147$ $0.276\%$
KuCoinPool $101$ $0.190\%$
EMCDPool $80$ $0.150\%$
PEGA Pool $76$ $0.143\%$
Table 3: The scale of mining hashrates of the most significant mining pools and independent miners of Bitcoin in 2022. Data sources: [1].
Figure 6: (a) Collective power of top mining pools in 2022 for selected Blockchains, monthly data [1][3][2]. (b) Collective number of hashes per second calculated by all miners in 2022 [$10^{20}$ hashes/second], weekly data [5].

### v2

@@ 7 Conclusions and Future Work
[…]
Appendix 0.A Discouraging Miners to Mine $txs_{1}$
Proof of the Theorem 4.4.
@@ Appendix 0.A Discouraging Miners to Mine $txs_{1}$
Let us denote with $\sigma_{1}$ all strategies where player $N_{s}$ continues to mine the chain with $txs_{1}$ in $\mathcal{S}_{p_{1},f_{1}}^{T}$, and let $\sigma_{2}$ denote all strategies, where he forks $txs_{1}$ in $\mathcal{S}_{p_{1},f_{1}}^{T}$ and then continues to mine the newly created chain (we assume that the transaction $txs_{1}$ was not mined by the miner $N_{s}$, as when $f_{2}-f>\frac{f_{1}-f}{\lambda_{s}}$, then as shown in the proof of the Theorem 4.3, the player $N_{s}$ will choose to mine $txs_{u}$ in $\mathcal{S}_{p_{1}}^{T}$). It is easy to see that once the player $N_{s}$ successfully forks $txs_{1}$ and creates a new chain, all miners other than the miner that mined $txs_{1}$ will prefer to mine the newly created chain. As the miner who mined the transaction $txs_{1}$ has the mining power less than $\lambda_{s}$, we can say that once $txs_{1}$ is forked, the miners with the mining power at least $1-\lambda_{s}$ will continue mining the new chain. In summary, for some constant $C$, given similar arguments as in the previous proofs:
@@
The strategy $\sigma_{2}$ is dominating for the player $N_{s}$ who did not mine $txs_{p_{1}}$, and he will prefer to continue mining $txs_{p_{1}}$ whenever $f_{2}-f>\frac{f_{1}-f}{\lambda_{s}}$. Note that the above analysis will be similar, even if some player other than the $N_{s}$ successfully starts a fork of $txs_{p_{1}}$.
@@
Appendix 0.B Average Hashrate
Figure 6: Average mining hashrate week by week of the most significant Bitcoin miners in 2022 [btcexplorer].